OpenStudy (atlchic):

solve by elimination,y-7=x and 2x-5y=-2

OpenStudy (atlchic):

show ur work

OpenStudy (karatechopper):

ah now i took an EOI exam today so u will be pleased to show u how;)

OpenStudy (atlchic):

k

OpenStudy (anonymous):

y-x=7 multiply both sides by 2 to get 2y-2x=14 -5y+2x=-2 now add -3y+0=12 y=-4 now plug -4 into either eq so -5(-4)+2x=-2 so 20+2x=-2 so 2x=-22 x=-11

OpenStudy (anonymous):

ok?

OpenStudy (atlchic):

answer my other question

OpenStudy (anonymous):

first of all, you should learn to ask, not demand. But where is the other question?

OpenStudy (atlchic):

plz i mean

OpenStudy (atlchic):

i sent u message

OpenStudy (anonymous):

y=5 x=2

OpenStudy (anonymous):

how many more you got? i am heading out

OpenStudy (atlchic):

um show ur work r work on dat problem plz

OpenStudy (atlchic):

3 more problems

OpenStudy (anonymous):

it is difficult to do a lot of problems for different people when your questions are in messages. please write out the question here. Thank you.

OpenStudy (atlchic):

k

OpenStudy (atlchic):

solve by elimination, 3x+0.2y=7 and 3x=0.4y+4. Show your work.

OpenStudy (atlchic):

hello

OpenStudy (anonymous):

3x+.2y=7 -3x+.4y=-4 so add and get .6=3 y= 5 now plug in 3x+.2(5)=7 3x+1=7 3x=6 x=2 do you always get people to do your homework?

OpenStudy (atlchic):

sorry like can you explain step by step, how to do elimination

hero (hero):

She wants you to literally, line-by-line, explain every step.

OpenStudy (atlchic):

no, can u show me how to do elimination, so i dont have 2 keep ang yall for answers

OpenStudy (anonymous):

you want to add the eqs in the following manner so that you eliminate one variable and therefore you can solve for th non-eliminated variable. you add as if it was columnar addition like 12 +34 ______ 46 so line up the eqs so that when you add like that, one variable will go to zero

hero (hero):

I can explain it, but only in vyew.

hero (hero):

This isn't the best platform for explaining elimination method.

OpenStudy (atlchic):

ok imma work an elimination problem on my own.

hero (hero):

Good luck

OpenStudy (atlchic):

x=y and x+y=7, solve by elimination.

OpenStudy (atlchic):

im lost can u help me